# The Negative Participation Paradox in Three-Candidate Instant Runoff Elections

### v1

@@ Abstract.
In this paper, we provide theoretical and empirical estimates for the likelihood of a negative participation paradox under instant runoff voting in three-candidate elections. We determine the probability of the paradox and related conditional probabilities based on the impartial anonymous culture and impartial culture models for both complete and partial ballots. We compare these results to the empirical likelihood of a negative participation paradox occurring under instant runoff voting in a large database of voter profiles which have been reduced to three candidates. Lastly, we analyze the relative likelihood of this paradox in comparison to other well-known paradoxes.
1. Introduction
@@
The voting method of instant runoff voting (IRV) is famously susceptible to many classical paradoxes. For example, under IRV an election can demonstrate an upward or downward monotonicity paradox [4, 16, 18, 22], or a no-show paradox [3, 4, 20]. IRV is also susceptible to paradoxes of negative participation, a lesser-studied type of paradox. The purpose of this article is to examine the likelihood that a three-candidate IRV election demonstrates this paradox.
@@
An election is said to demonstrate a negative participation paradox under a given voting method if the addition of voters who rank candidate $A$ last causes $A$ to go from losing to winning. This paradox is closely related to the notion of a positive participation paradox [11], also referred to in the language of “positive involvement” [25], where the addition of voters who rank candidate $A$ uniquely first causes $A$ to go from winning to losing. Such outcomes are sometimes referred to as strong no-show paradoxes [2, 24]. To the best of our knowledge, positive participation paradoxes have received more attention in prior literature than negative participation paradoxes, even though IRV is not susceptible to the positive type. Prior research on positive participation paradoxes focuses on other voting methods such as Condorcet methods [2, 8, 9, 24] (this literature tends to use the term “strong no-show paradox” instead of positive participation paradox). Much of the prior work concerning negative participation paradoxes also examines the susceptibility of Condorcet methods to this paradox [2, 24], while some focuses on scoring runoff rules like IRV [11, 13]. Much previous work on these kinds of voting paradoxes focuses on elections with a small number of candidates, often analyzing elections with only three candidates. Our work similarly focuses on the three-candidate case.
@@
This paper makes one small and two substantive contributions regarding the likelihood that an election demonstrates a negative participation paradox under IRV in a three-candidate election. First, for our small contribution we provide new theoretical results under the impartial anonymous culture (IAC) model of voter behavior. Most of the interesting work in this vein has been done in [11, 13], and our results in this area fill some small gaps left from those papers. Second, we provide results under the impartial culture (IC) model of voter behavior. We are able to obtain these new results using geometric techniques developed in [10] and [26]. Third, we provide empirical results using a large database of IRV elections from American and Scottish political elections, as well as elections from the American Psychological Association. Ours is the first study which provides such empirical results; the only other previous work which searched for negative participation paradoxes in real-world ballot data is [15], which focuses on multiwinner single transferable vote elections from Scotland. Our general finding is that negative participation paradoxes occur at a much higher rate than other paradoxes.
@@
Because we are interested in providing both theoretical and empirical analyses of the negative participation paradox we consider situations when all voters provide complete rankings and when voters are allowed to provide only partial ballots. This allows for a better comparison between theoretical and empirical results since partial ballots are extremely common in real-world election data. On the theoretical side, we provide estimates that the likelihood of a negative participation paradox occurs under IRV for 3 candidates under under both the IAC model and IC models. For the IAC model, we analyze the complete and partial ballot cases using the software package Normaliz and Monte Carlo simulations, respectively. For the IC model we use geometric techniques for both full and partial ballots.
@@
The widespread use of the method of single transferable vote in Scotland and other locations, as well as the recent rise of IRV in the US, has made available a large collection of ranked data from real elections. When an election contained more than three candidates, we reduce each election to three candidates by running a number of initial rounds of IRV. Then we determine the number of elections susceptible to a negative participation paradox using the existing ballot data as well as data which we make more “complete” using numerical techniques which we describe in Section 4.
@@
2. Preliminaries
Instant-runoff voting (IRV), often colloquially referred to as “ranked-choice voting”, has become increasingly popular in the United States as an alternative to the method of plurality when there are more than 2 candidates. In such elections voters cast preference ballots with a (possibly partial) linear ranking of the candidates. After an election, the ballots are aggregated into a preference profile, which shows the number of each type of ballot cast. For example, Table 1 shows a preference profile involving the three candidates $J$, $S$, and $W$. The number 1430 denotes that 1430 voters ranked $J$ first, $S$ second, and $W$ third. We use $\succ$ to denote when a voter ranks one candidate over another, so that 1430 voters cast the ballot $J\succ S\succ W$. An IRV election proceeds in rounds. In each round, the number of first-place votes for each candidate is calculated. If a candidate has a majority of the first-place votes, they are declared the winner. If no candidate receives a majority of votes, the candidate with the fewest first-place votes is eliminated and the names of the remaining candidates on the affected ballots are moved up. The process repeats until a candidate is declared the winner [^1] The issue of ties does not arise in our work in this paper, and thus we ignore the possibility of multiple winners.. In the case when partial ballots are allowed, any ballot in which all candidates have been eliminated is considered “exhausted” and plays no further role in the determination of the winner.
When there are only three candidates, at most 2 rounds are required to determine a winner; in general, several rounds may be required. In our analysis of the real-world database in Section 4, we run a sufficient number of rounds for each set of voter preferences until the number of candidates is reduced to three and consider the resulting three-candidate election. We demonstrate IRV in such a case in the following example.
@@ 2. Preliminaries
Table 1. Vote totals for the top 3 candidates in the 2023 election for Minneapolis City Council Ward 8
It is well-known that IRV can produce strange outcomes when ballots are added or removed from an election [3, 4, 5, 6, 11, 13, 15, 16], or when candidates are added or removed from an election [7, 17, 19, 28]. The example below illustrates a type of paradoxical outcome which can occur when ballots are added to an election.
@@
Suppose we add 4000 ballots of the form $W\succ J\succ S$ to the ballots from in Table 1. What effect should this have on the electoral outcome? Since these voters all prefer $W$, it would make sense for White to win after receiving this large boost of support. Since these voters prefer Jenkins to Stevenson, it would also make sense for these ballots to bolster Jenkins’ victory. However, the addition of these ballots causes Jenkins to be eliminated before White and Stevenson edges out White in the final round.
Thus, adding thousands of ballots in which Stevenson is ranked last can result in a good outcome for him. Note that this outcome is also true in the original four-candidate election including Sullentrop ($Su)$: the same result is obtained if we add 4000 ballots of the form $W\succ J\succ Su\succ S$ or 4000 ballots of the form $W\succ Su\succ J\succ S$. That is, when we add ballots to create a negative participation paradox in an IRV election which has been reduced to three candidates, only the first and last rankings matter. We can fill in the intermediate rankings in an arbitrary fashion; the original election with all candidates also demonstrates a negative participation paradox as long as sufficient votes are added so that the plurality loser remains until eliminated in the final round.
@@
That is, an election demonstrates this paradox when a losing candidate can be made into a winner by adding ballots on which the loser is ranked last. We use the language of “negative participation paradox” following [11], but there is no standard vocabulary around the topic. For example, much of the literature concerning these paradoxes and Condorcet methods [8, 9, 24] uses the the language of no-show paradoxes and negative and positive involvement. The previous example shows that IRV is susceptible to negative participation paradoxes; we note that other voting methods such as positional scoring rules cannot exhibit this type of paradox.
@@
In a three-candidate election, a negative participation paradox can only occur when several conditions are satisfied. First, the added votes cannot result in the plurality loser being made into the IRV winner: if the added votes rank the plurality loser last, the candidate will remain the plurality loser and be eliminated in the first round. Second, the IRV winner must be different from the plurality winner. To see this, suppose that the IRV winner is $A$ and the candidates are ranked by first-place votes $A>B>C.$ Since $A$ is the IRV winner, $B$ must be the losing candidate who can be made into a winner by the addition of votes ranking $B$ last. But $B$ cannot be the winner in such a fashion: if $B$ becomes the new plurality loser then they are eliminated in the first round, and if $C$ remains the plurality loser than $C$ is eliminated in the first round, leaving $B$ to be eliminated in the second round as they were in the original election (the added votes increase the margin of $A$ over $B$). Hence the IRV and plurality winners must be distinct, implying it is the plurality winner who can be made to win under IRV by the addition of votes ranking them last. Similar reasoning implies that the new votes must rank the plurality loser first so that the original IRV winner is eliminated in the first round of the new election.
We use this logic to determine the probability of a negative participation paradox when there are three candidates in the next section.
3. The Impartial Anonymous and Impartial Culture Models
@@
Table 2. Preference profile with complete preferences.
@@ C B C A B A
Table 3. Preference profile with some partial preferences.
@@
In the complete ballot case, we can assume WLOG that
(1) $a_{1}+a_{2}>b_{1}+b_{2}>c_{1}+c_{2};$
similarly, in the partial ballot case we assume WLOG that
(2) $a_{0}+a_{1}+a_{2}>b_{0}+b_{1}+b_{2}>c_{0}+c_{1}+c_{2}.$
That is, we assume $A$ is the plurality winner and $C$ is the plurality loser. As described in Section 2, a negative participation paradox can only occur when the plurality ($A$) and IRV ($B$) winners are distinct and when additional votes are added with the ranking $C\succ B\succ A.$ Hence, initially
(3) $a_{1}+a_{2}+c_{1}<b_{1}+b_{2}+c_{2}.$
@@
(4) $b_{1}+b_{2}-(c_{1}+c_{2})<a_{1}+a_{2}+b_{1}-(b_{2}+c_{1}+c_{2})\quad\mbox{ or }\quad a_{1}+a_{2}>2b_{2}.$
It is easily seen that (3) and (4) are also sufficient to ensure a negative participation paradox can occur given that $A$ is the plurality winner and $C$ is the plurality loser.
@@
Before defining and analyzing the two models, we note that they are widely used in the social choice literature to provide a priori estimates for the probabilities of many phenomena in social choice, and they tend to provide theoretical upper bounds for such probabilities. For example, in three-candidate elections the probability that an election fails to contain Condorcet winner is 6.25% under the IAC model (this is easily calculated using Normaliz); the probability falls essentially to 0 in real-world elections [6, 21, 27]. Thus, the models are useful for providing theoretical “worst-case” estimates for various election probabilities.
3.1. Impartial Anonymous Culture
@@
Under the impartial anonymous culture (IAC) model, the vector of vote totals at the top of the preference profile is assumed to be chosen at random. For example, in the complete ballot case we choose at random a vector of the form $(a_{1},a_{2},b_{1},b_{2},c_{1},c_{2})$ subject to the constraints that each entry is non-negative and the six vote totals sum to the number of voters. Since we examine only limiting probabilities, we choose a vector of proportions of length six in the complete ballot case or a vector of proportions of length nine in the partial ballot case. Limiting probabilities involving IAC can be determined using the theory of Ehrhart polynomials, which has been implemented in the software package Normaliz [1]. The probabilities presented in Proposition 1 were calculated using Normaliz. The probability of a negative participation paradox using IRV assuming complete rankings appears in [11]. In the following proposition, we include this result along with related conditional probabilities for reference.
@@
Assuming, as in the previous section, that first-place vote ranking is $A>B>C,$ the probability that an election demonstrates a negative participation paradox is equal to the probability that (3) and (4) are satisfied. The results in (ii) and (iii) below are conditional probabilities. If an election demonstrates a negative participation paradox then there cannot be a majority candidate (since the plurality and IRV winners would coincide). Thus, in (ii), the probability that an election demonstrates a negative participation paradox winner assuming the election does not contain a majority candidate is equal to the the probability of a negative participation paradox divided by the probability that an election does not have a majority candidate. Likewise, in (iii), the probability that an election demonstrates a negative participation paradox assuming the IRV winner is not the plurality winner is equal to the probability that an election demonstrates a negative participation paradox divided by the probability the IRV winner is not the plurality winner.
@@
    the limiting probability that an election demonstrates a negative participation paradox is $7/96=7.29\%$ [11].
@@
    the limiting probability that an election demonstrates a negative participation paradox is $7/54=12.96\%$, assuming the election does not contain a majority candidate.
@@
We are unaware of analytical methods which extend the Ehrhart polynomial techniques employed by Normaliz to the partial ballot setting where a preference profile has nine parameters. Thus, to estimate the corresponding IAC probabilities under partial ballots we use Monte Carlo simulation. For each run, we choose a percentage distribution of length nine at random (straightforward techniques exist for this task, see [12] for example) and check for a paradox. Using 100,000 runs, in the partial ballot setting we estimate the corresponding limiting probabilities from Proposition 1 to be approximately 7.24%, 16.74%, and 62.49%.
3.2. Impartial Culture
Under the impartial Culture (IC) model, each voter selects uniformly and independently among all possible rankings of the candidates. We can approximate the likelihood of a negative participation paradox for a large number of voters by calculating the limiting probability as $V\rightarrow\infty$ using the geometric methods introduced in [26] and subsequently elaborated on in [10]. Note that the probability that there is a majority candidate is zero in the limiting case, thus we provide counterpoints to only two of the probabilities identified in Proposition 1.

### v3

@@ Abstract
We provide theoretical and empirical estimates for the likelihood of a negative participation paradox under instant runoff voting in three-candidate elections. We determine the probability of the paradox and related conditional probabilities based on the impartial anonymous culture and impartial culture models for both complete and partial ballots. We compare these results to the empirical likelihood of a negative participation paradox occurring under instant runoff voting in a large database of voter profiles which have been reduced to three candidates. Lastly, we analyze the relative likelihood of this paradox in comparison to other well-known paradoxes.
keywords:
negative participation paradox, negative involvement, instant runoff voting, impartial culture, empirical results, voting paradox
1 Introduction
@@
The voting method of instant runoff voting (IRV) is famously susceptible to many classical paradoxes. For example, under IRV an election can demonstrate an upward or downward monotonicity paradox (Fishburn & Brams, 1983; McCune and Graham-Squire, 2023; Miller, 2017; Ornstein & Norman, 2014), or a no-show paradox (Felsenthal & Tideman, 2013; Fishburn & Brams, 1983; Moulin, 1988). IRV is also susceptible to paradoxes of negative participation, a lesser-studied type of paradox. The purpose of this article is to examine the likelihood that a three-candidate IRV election demonstrates this paradox.
@@
An election is said to demonstrate a negative participation paradox under a given voting method if the addition of voters who rank candidate $A$ last causes $A$ to go from losing to winning. We use the term “negative participation paradox” following (Kamwa et al., 2023), but other authors use language such as “violation of negative involvement” (Pérez, 2001). This paradox is closely related to the notion of a positive participation paradox (Kamwa et al., 2023), also referred to in the language of “positive involvement” (Saari, 1995), where the addition of voters who rank candidate $A$ uniquely first causes $A$ to go from winning to losing. Such outcomes are sometimes referred to as strong no-show paradoxes (Duddy, 2014; Pérez, 2001). To the best of our knowledge, positive participation paradoxes have received more attention in prior literature than negative participation paradoxes, even though IRV is not susceptible to the positive type. Prior research on positive participation paradoxes focuses on other voting methods such as Condorcet methods (Duddy, 2014; Holliday, 2024; Holliday & Pacuit, 2021; Pérez, 2001); this literature tends to use the term “strong no-show paradox” instead of positive participation paradox. Much of the prior work concerning negative participation paradoxes also examines the susceptibility of Condorcet methods to this paradox (Duddy, 2014; Pérez 2001), while some focuses on scoring runoff rules like IRV (Kamwa et al., 2023, Lepelley & Merlin, 2001). Much previous work on these kinds of voting paradoxes focuses on elections with a small number of candidates, often analyzing elections with only three candidates. Our work similarly focuses on the three-candidate case.
@@
This paper makes one small and two substantive contributions regarding the likelihood that an election demonstrates a negative participation paradox under IRV in a three-candidate election. First, for our small contribution we provide new theoretical results under the impartial anonymous culture (IAC) model of voter behavior. Most of the interesting work in this vein has been done in (Kamwa et al., 2023, Lepelley & Merlin, 2001), and our results in this area fill some small gaps left from those papers. Second, we provide results under the impartial culture (IC) model of voter behavior. We are able to obtain these new results using geometric techniques developed in (Kamwa & Merlin, 2015) and (Saari & Tataru, 1999). Third, we provide empirical results using a large database of IRV elections from American and Scottish political elections, as well as elections from the American Psychological Association. Ours is the first study which provides such empirical results; the only other previous work which searched for negative participation paradoxes in real-world ballot data is (McCune, 2024), which focuses on multiwinner single transferable vote elections from Scotland. Our general finding is that negative participation paradoxes occur at a much higher rate than other paradoxes.
@@
Because we are interested in providing both theoretical and empirical analyses of the negative participation paradox we consider situations when all voters provide complete rankings and when voters are allowed to provide only partial rankings. This allows for a better comparison between theoretical and empirical results since partial ballots are extremely common in real-world election data. On the theoretical side, we provide estimates that the likelihood of a negative participation paradox occurs under IRV for three candidates under under both the impartial anonymous culture (IAC) and impartial culture (IC) models. For the IAC model, we analyze the complete and partial ballot cases using the software package Normaliz and Monte Carlo simulations, respectively. For the IC model we use geometric techniques for both full and partial ballots.
@@
The widespread use of the method of single transferable vote in Scotland and other locations, as well as the recent increase in the use of IRV in the US, has made available a large collection of preference ballot data from real-world elections. When an election contained more than three candidates, we reduce each election to three candidates by running a number of initial rounds of IRV. We then determine the number of elections susceptible to a negative participation paradox using the existing ballot data as well as data which we make more “complete” using numerical techniques which we describe in Section 4.
@@
2 Preliminaries
Instant-runoff voting (IRV), often colloquially referred to as “ranked-choice voting”, has become increasingly popular in the United States as an alternative to the method of plurality. In an election which uses IRV voters cast preference ballots with a (possibly partial) linear ranking of the candidates. After an election, the ballots are aggregated into a preference profile, which shows the number of each type of ballot cast. For example, Table 1 shows a preference profile involving the three candidates $J$, $S$, and $W$. The number 1430 denotes that 1430 voters ranked $J$ first, $S$ second, and $W$ third. We use $\succ$ to denote when a voter ranks one candidate over another, so that 1430 voters cast the ballot $J\succ S\succ W$. An IRV election proceeds in rounds. In each round, the number of first-place votes for each candidate is calculated. If a candidate has a majority of the (remaining) first-place votes, they are declared the winner. If no candidate receives a majority of votes, the candidate with the fewest first-place votes is eliminated and the names of the remaining candidates on the affected ballots are moved up. The process repeats until a candidate is declared the winner. [^1] The issue of ties does not arise in our work in this paper, and thus we ignore the possibility of multiple winners. In the case when partial ballots are allowed, any ballot in which all candidates have been eliminated is considered “exhausted” and plays no further role in the determination of the winner.
When there are only three candidates, at most two rounds are required to determine a winner; in general, several rounds may be required. In our analysis of the real-world database in Section 4, we run a sufficient number of rounds for each set of voter preferences until the number of candidates is reduced to three and consider the resulting three-candidate election. We demonstrate IRV in such a case in the following example.
@@ 2 Preliminaries
Table 1: Vote totals for the top 3 candidates in the 2023 election for Minneapolis City Council Ward 8
It is well-known that IRV can produce strange outcomes when ballots are added or removed from an election (Felsenthal & Tideman, 2013; Fishburn & Brams, 1983; Graham-Squire, 2024; Graham-Squire and McCune, 2023; Kamwa et al., 2023; Lepelley & Merlin, 2001; McCune & Graham-Squire, 2023), or when candidates are added or removed from an election (Green-Armytage et al., 2016; McCune & Wilson, 2023; Miller, 2019; Tideman, 1987). The example below illustrates a type of paradoxical outcome which can occur when ballots are added to an election.
@@
Suppose we add 3000 ballots of the form $W\succ J\succ S$ to the ballots from in Table 1. What effect should this have on the electoral outcome? Since these voters all prefer $W$, it would make sense for White to win after receiving this large boost of support. Since these voters prefer Jenkins to Stevenson, it would also make sense for these ballots to bolster Jenkins’ victory. However, the addition of these ballots turns Stevenson into the winner. To see this, note that with these ballots added to the electorate the initial first-place vote totals are 3591, 3640, and 3668 for Jenkins, Stevenson, and White, respectively. In this modified election Jenkins is eliminated first and after the vote transfers Stevenson defeats White 5070 votes to 4680.
Thus, adding thousands of ballots in which Stevenson is ranked last can result in a good outcome for him. Note that this outcome is also true in the original four-candidate election including Sullentrop ($Su)$: the same result is obtained if we add 3000 ballots of the form $W\succ J\succ Su\succ S$ or 3000 ballots of the form $W\succ Su\succ J\succ S$. That is, when we add ballots to create a negative participation paradox in an IRV election which has been reduced to three candidates, only the first and last rankings matter. We can fill in the intermediate rankings in an arbitrary fashion; the original election with all candidates also demonstrates a negative participation paradox as long as sufficient votes are added so that the original three-candidate plurality loser remains until eliminated in the penultimate round.
@@
That is, an election demonstrates this paradox when a losing candidate can be made into a winner by adding ballots on which the loser is ranked last. We use the language of “negative participation paradox” following (Kamwa et al., 2023), but, as noted in the Introduction, there is no standard vocabulary around the topic. For example, much of the literature concerning these paradoxes and Condorcet methods (Holliday, 2024; Holliday & Pacuit, 2021; Pérez, 2001) uses the the language of strong no-show paradoxes and negative and positive involvement. We note that for some researchers, a negative participation paradox can be considered a special case of a no-show paradox, a paradox briefly discussed in Section 5. This is not the case for us, as we use the definition of no-show paradox (sometimes also referred to as an abstention paradox) found in (Graham-Squire, 2024) and (Kamwa et al., 2023). In these articles, a three-candidate election demonstrates a no-show paradox if there exists a set of ballots $\mathcal{B}$ with the election winner $A$ ranked last such that if we remove these ballots then $A$ loses in the resulting election.
Example 2 shows that IRV is susceptible to negative participation paradoxes; many other voting methods such as positional scoring rules cannot exhibit this type of paradox. We also note that other researchers have studied the notion of a positive participation paradox, which occurs if there exists a winning candidate $A$ and a set of ballots $\mathcal{B}$ ranked first such that the addition of these ballots to the electorate turns $A$ into a losing candidate. We do not analyze positive participation paradoxes in this article because IRV is not susceptible to it; however, the voting method of single transferable vote, which is the multiwinner version of IRV, is susceptible to positive participation paradoxes (see (McCune, 2024) for two real-world examples).
@@
In a three-candidate election, a negative participation paradox can occur only when several conditions are satisfied. First, the added votes cannot result in the plurality loser being made into the IRV winner: if the added votes rank the plurality loser last, that candidate will remain the plurality loser and be eliminated in the first round. Second, the IRV winner must be different from the plurality winner. To see this, suppose that the IRV winner is $A$ and the candidates are ranked by first-place votes $A>B>C.$ Since $A$ is the IRV winner, $B$ must be the losing candidate who can be made into a winner by the addition of votes ranking $B$ last. But $B$ cannot be the winner in such a fashion: if $B$ becomes the new plurality loser then they are eliminated in the first round, and if $C$ remains the plurality loser than $C$ is eliminated in the first round, leaving $B$ to be eliminated in the second round as they were in the original election (the added votes increase the margin of $A$ over $B$). Hence the IRV and plurality winners must be distinct, implying it is the plurality winner who can be made to win under IRV by the addition of votes ranking them last. Similar reasoning implies that the new votes must rank the plurality loser first so that the original IRV winner is eliminated in the first round of the new election.
We use this logic to determine the probability of a negative participation paradox in three-candidate elections under two models of voter behavior in the next section.
3 The Impartial Anonymous and Impartial Culture Models
@@ C B C A B A
Table 2: Preference profile with complete preferences.
@@
Table 3: Preference profile with some partial preferences.
In the complete ballot case, we can assume without loss of generality that
$a_{1}+a_{2}>b_{1}+b_{2}>c_{1}+c_{2};$ (1)
similarly, in the partial ballot case we assume without loss of generality that
$a_{0}+a_{1}+a_{2}>b_{0}+b_{1}+b_{2}>c_{0}+c_{1}+c_{2}.$ (2)
That is, we assume $A$ is the unique plurality winner and $C$ is the unique plurality loser. As described in Section 2, a negative participation paradox can only occur when the plurality ($A$) and IRV ($B$) winners are distinct and when additional votes are added with the ranking $C\succ B\succ A.$ Hence, initially
$a_{1}+a_{2}+c_{1}<b_{1}+b_{2}+c_{2}.$ (3)
@@
$b_{1}+b_{2}-(c_{1}+c_{2})<a_{1}+a_{2}+b_{1}-(b_{2}+c_{1}+c_{2})\quad\mbox{ or }\quad a_{1}+a_{2}>2b_{2}.$ (5)
It is easily seen that (3) and (5) are also sufficient to ensure a negative participation paradox can occur given that $A$ is the plurality winner and $C$ is the plurality loser [^2] In the non-limiting case where the number of voters is finite, (4) must have room for at least one integer value of $z$ to fit in the interval, and thus in this case (5) becomes $a_{1}+a_{2}>2b_{2}+1$.. Moreover, (4) provides an explicit formula for the range of added votes that result in the paradox.
@@
Before defining and analyzing the two models, we note that they are widely used in the social choice literature to provide a priori estimates for the probabilities of many phenomena in social choice, and they tend to provide theoretical upper bounds for such probabilities. For example, in three-candidate elections the probability that an election fails to contain Condorcet winner is 6.25% under the IAC model (this is easily calculated using Normaliz); the probability falls essentially to 0 in real-world elections (Graham-Squire & McCune, 2023; Myers, 2024; Song, 2023). Thus, the models are useful for providing theoretical “worst-case” estimates for various election probabilities.
3.1 Impartial Anonymous Culture
@@
Under the impartial anonymous culture (IAC) model, the vector of vote totals at the top of the preference profile is assumed to be chosen at random. For example, in the complete ballot case we choose at random a vector of the form $(a_{1},a_{2},b_{1},b_{2},c_{1},c_{2})$ subject to the constraints that each entry is non-negative and the six vote totals sum to the number of voters. Since we examine only limiting probabilities, we choose a vector of proportions of length six in the complete ballot case or a vector of proportions of length nine in the partial ballot case. Limiting probabilities involving IAC can be determined using the theory of Ehrhart polynomials, which has been implemented in the software package Normaliz (Bruns et al., 2022). The probabilities presented in Proposition 1 were calculated using Normaliz. The probability of a negative participation paradox using IRV assuming complete rankings appears in (Kamwa et al., 2023). In the following proposition, we include this result along with related conditional probabilities for reference.
@@
Assuming, as in the previous section, that the first-place vote ranking is $A>B>C,$ the probability that an election demonstrates a negative participation paradox is equal to the probability that (3) and (5) are satisfied. This probability, shown in (i), is obtained directly from (Kamwa et al., 2023). The results in (ii) and (iii) below are conditional probabilities, obtained directly from Normaliz. If an election demonstrates a negative participation paradox then there cannot be a majority candidate (a candidate who receives a majority of first-place votes) because the plurality and IRV winners would coincide. Thus, in (ii), the probability that an election demonstrates a negative participation paradox winner assuming the election does not contain a majority candidate is equal to the the probability of a negative participation paradox divided by the probability that an election does not have a majority candidate. Likewise, in (iii), the probability that an election demonstrates a negative participation paradox assuming the IRV winner is not the plurality winner is equal to the probability that an election demonstrates a negative participation paradox divided by the probability the IRV winner is not the plurality winner. Previous studies such as (Miller, 2017) have shown that as an election gets “closer” the probability of observing various voting paradoxes increases; items (ii) and (iii) of the proposition also demonstrate this for negative participation paradoxes.
@@
    the limiting probability that an election demonstrates a negative participation paradox is $7/96=7.29\%$ (Kamwa et al., 2023).
@@
    the limiting probability that an election demonstrates a negative participation paradox is $1/6=16.67\%$, assuming the election does not contain a majority candidate.
@@
Normaliz also allows us to obtain the liminting probabilities under IAC for the partial ballot case.
Proposition 2.
In a 3-candidate election in which voters can provide partial preferences, under IAC
  1. i
    the limiting probability that an election demonstrates a negative participation paradox is $169/2304=7.34\%$.
  2. ii
    the limiting probability that an election demonstrates a negative participation paradox is $169/1305=12.95\%$, assuming the election does not contain a majority candidate.
  3. iii
    the limiting probability that an election demonstrates a negative participation paradox is $117/187=62.57\%$, assuming the IRV winner is not the plurality winner.
3.2 Impartial Culture
Under the impartial Culture (IC) model, each voter selects uniformly and independently among all possible rankings of the candidates. We can approximate the likelihood of a negative participation paradox for a large number of voters by calculating the limiting probability as $V\rightarrow\infty$ using the geometric methods introduced in (Saari & Tataru, 1999) and subsequently elaborated on in (Kamwa & Merlin, 2015). The probability that there is a majority candidate is zero in the limiting case under the IC model, and thus we provide counterpoints to only two of the probabilities identified in Proposition 1.